You recently financed the purchase of a new home with a $200,000 mortgage at a 7.5% annual interest rate over 30 year. They want to review how they are doing after 18 months and have asked you to assist then with answers to the following questions: How much will they have paid on their original mortgage balance after 18 months? How much total interest will they have paid at that point? What will be their unpaid principal balance on the mortgage at that point
Answer #1

Calculating Monthly Payment on Loan,

using TVM Calculation,

PMT = [PV = 200,000, FV = 0 N = 360, I = 0.075/12]

PMT = $1,398.43

a.

Calculating Mortgage Balance after 18 months,

FV = [PV = 200,000, PMT = -1,398.43, N = 18, I = 0.075/12]

FV = $197,181.48

Mortgage Paid = (200,000 - 197,181.48) = $2,818.52

b.

Interest Paid = 18(1,398.43) - (200,000 - 197,181.48)

Interest Paid = $25,171.74

c.

Unpaid Balance = $197,181.48